# Rank jumps and Multisections of elliptic fibrations on K3 surfaces

### v1

@@ 1. Introduction
Let $X$ be a K3 surface defined over a number field such that $\operatorname{Pic}(X)=U\oplus\langle-2d\rangle$, for $d\in\mathbb{N}$, or equivalently a K3 surface admitting an elliptic fibration and with Picard number 3. Assume that $d\neq 2,3$. Then each elliptic fibration on $X$ has the potential rank jump property and, moreover, if $d\equiv 3\mod 4$, then it has the rank jump property over its field of definition.
@@
    The generic member in the 18-dimensional family of $U(2)$–polarized K3 surfaces admits a saliently ramified smooth bisection of genus 1 (Proposition 3.6).
@@ 2.1. Genus 1 fibrations and Elliptic surfaces
The set of $k$-sections of $\pi$ is an abelian group which we denote by $\operatorname{MW}(X,\pi,k)$. We often omitt $k$, if the field of definition is clear from the context. The fibration $\pi:X\to C$ is called trivial if its generic fiber is isomorphic to $E\otimes_{k}k(C)$ for some elliptic curve $E$ over $k$.
@@ 2.1.1. Multisections of genus 1 fibrations
Given a genus 1 fibration $\pi:X\rightarrow C$, a geometrically integral curve $M\subset X$ is called a multisection of $\pi$ if $\pi|_{M}:M\rightarrow C$ is finite flat and of degree larger than 1.
@@
Our definition is slightly different from that in [BT00, Def. 3.2.]. We assume that the degree of the restriction of the fibration is larger than 1, while [BT00] assume that it is non-zero. So, in our definition, sections are not multisections. This modification is natural in our settings. Indeed, we are interested in the ramification of the restriction of the fibration of the multisection (Def. 2.5) which clearly only makes sense for maps of degree at least 2.
@@ 2.2. Rank jump and base change
Let $M\subset X$ be a multisection of $\pi$ defined over $l/k$. Let $X_{M}$ be the surface obtained by resolving the singularities of $X\times_{C}M$. The projection onto $M$ induces an elliptic fibration $\pi_{M}:X_{M}\rightarrow M$ and we have an injection $\alpha:\mathrm{MW}(X,\pi,l)\hookrightarrow\mathrm{MW}(X_{M},\pi_{M},l)$ yielding
@@
Let $\iota:M\rightarrow X$ be the inclusion map and $\mathrm{id:}M\rightarrow M$, the identity map. Then $\sigma_{M}=(\iota,\mathrm{id}):M\rightarrow X_{M}$ is a section of $\pi_{M}$. If the latter is linearly independent of the sections in the image of $\alpha$ then the inequality in (2.1) is strict and, in particular, an application of Silverman’s Specialization Theorem ([Sil83, Theorem C]) implies that $\mathrm{rank}X_{t}(l)>\mathrm{rank}\,\operatorname{MW}(X,\pi,l)$ for every $t$ in the image of the $l$-points of the map $\pi|_{M}:M\rightarrow C$.
@@
Let $\pi:X\rightarrow C$ be an elliptic fibration and $M$ a saliently ramified multisection defined over a field $l$. Let $X_{M}$ be the resolution of $X\times_{C}M$ and $\pi_{M}:X_{M}\rightarrow M$ be the elliptic fibration induced by $\pi$. Then the section induced on $X_{M}$ by $\sigma_{M}=(\iota,\mathrm{id}):M\rightarrow X\times_{C}M$ is linearly independent of the pull-back of the sections of $\pi$ in $\operatorname{MW}(X_{M},\pi_{M},l)$.
@@
$\mathcal{R}(X,\pi,l)\text{ is infinite.}$.
@@ 2.3. Lattice polarized K3 surfaces and genus 1 fibrations
([Ko, Lemma 2.1]) A K3 surface $X$ admits an elliptic fibration if, and only if, the lattice $U\simeq\left[\begin{array}[]{ll}0&1\\ 1&0\end{array}\right]$ is primitively embedded in $\operatorname{Pic}(X)$. In particular, the family of K3 surfaces admitting an elliptic fibration is 18-dimensional and coincides with the family of $U$-polarized K3 surfaces.
@@
This result is well known, cf. [Ko, Lemma 2.1]. We briefly recall the idea of the proof. If a K3 surface admits an elliptic fibration, then there exists a smooth genus 1 curve, that is, a fiber of the fibration. We denote it by $F$ and observe that $F^{2}=0$ (by adjuction). Moreover, there exists a rational curve that is a section, denoted by $\mathcal{O}$. Then $\mathcal{O}^{2}=-2$ (by adjuction) and $F\mathcal{O}=1$ (since $\mathcal{O}$ is a section). Therefore, the lattice $U\simeq\langle F,F+\mathcal{O}\rangle$ is a sublattice of $\operatorname{Pic}(X)$. Since $U$ is a unimodular lattice, it does not have overlattices, so the embedding $\langle F,F+\mathcal{O}\rangle\subset\operatorname{Pic}(X)$ is primitive. The converse is more delicate; indeed, it is not true that each class with self-intersection 0 is necessarily the class of a smooth irreducible curve (and hence of a fiber of a genus 1 fibration). Nevertheless, given a primitive class in $\operatorname{Pic}(X)$ with self-intersection 0, it is true that up to reflections induced by smooth rational curves on the surface, one can assume that it is the class of a smooth irreducible curve, which necessarily has genus 1 by adjuction. So, let $U$ be primitively embedded in $\operatorname{Pic}(X)$. Considering (potentially) reflections in $-2$-curves, one can assume that one of the classes with self-intersection 0 is the class of a smooth irreducible curve of genus 1, say $F$, and then $|F|$ defines a fibration onto $\mathbb{P}(H^{0}(X,\mathcal{O}_{X}(F))^{\vee})=\mathbb{P}^{1}$. Denoted $\{F,v\}$ the basis of $U$, with $F$ as above, $v-F$ is a $-2$-class with positive intersection with $F$, i.e., it is an effective $-2$-class. Moreover, it intersects the fiber in one point and therefore it is either an irreducible curve, and in particular a section, or the union of several curves, among which there is a section. In both the cases the fibration admits a section.∎
@@
If a K3 surface contains two distinct genus 1 fibrations, then the class of the fiber of the first one, say $F_{1}$, and the one of the second one, say $F_{2}$, have a positive intersection $F_{1}F_{2}=m>1$. The intersection has to be larger than 1 since the basis of a genus 1 fibration on a K3 surface is rational and so $F_{2}$ cannot be a section. In particular, $\langle F_{1},F_{2}\rangle\simeq U(m)$, with $m>1$, is primitively embedded in the Picard group of the surface. The generic K3 surface $X$ such that $\operatorname{Pic}(X)\simeq U(m)$ admits two genus 1 fibrations whose classes of fibers meet in $m$ points. Therefore the K3 surfaces that admit more than one genus 1 fibration are contained in the families of $U(m)$-polarized K3 surfaces for a certain $m>1$.
@@ 2.3.3. K3 surfaces with an elliptic fibration and low Picard number
Let $(X,\pi,\mathbb{P}^{1})$ be an elliptic fibration on the K3 surface $X$. Then there are no multiple fibers of $\pi$, due to the presence of a section, nevertheless there can be reducible fibers which contain some non-reduced components. These are the fibers of type $I_{n}^{*}$, $II^{*}$, $III^{*}$ or $IV^{*}$ in the Kodaira classification of the singular fibers, see [Mir89] for notation. Each of these fibers has at least 5 components, and in particular at least 4 components which do not intersect the zero section. These components are independent in $\operatorname{Pic}(X)$, and they are also independent with respect to the classes $F$ and $\mathcal{O}$ of the fiber and of the zero section of $\pi$, respectively. Hence, the presence of a fiber with non-reduced components forces the Picard number of the surface to be at least 6.
@@
Let $X$ be a K3 surface with $\rho(X)\leq 6$. Any elliptic fibration $\pi:X\rightarrow\mathbb{P}^{1}$ admits neither fibers with non-reduced components nor torsion sections.
@@ 3. Bisections of genus 1 and elliptic fibrations
For any other $d$, any elliptic fibration on $X$ has no reducible fibers, the Mordell–Weil group of any elliptic fibration on $X$ is isomorphic to $\mathbb{Z}$ and admits a generator that intersects the zero section in $d-2$ points.
@@ 3.1. Bisections of genus 1 fibrations
Let $X$ be a K3 surface that admits a genus 1 fibration $\pi:X\rightarrow\mathbb{P}^{1}$. The class of the fiber $F$ has trivial self-intersection, and there exists an ample class $H$ with $H^{2}=d>0$ among the generators of $\operatorname{Pic}(X)$, whose rank is forced to be at least 2. In other words, a general K3 surface $X$ with a genus 1 fibration and Picard number 2 has $\operatorname{Pic}(X)\simeq\Gamma_{b,c}$, where $\Gamma_{b,c}$ is the indefinite lattice of rank 2 defined as
@@
where one can assume $0\leq c<b$, see [vG05, Remark 4.2].
@@
If $X$ admits a genus 1 fibration with a bisection and $\rho(X)=2$, then $\operatorname{Pic}(X)=\Gamma_{2,c}$ and by [vG05, Proposition 3.7], we can assume that $c=0$ or $c=1$. Let $D_{1},D_{2}$ be two generators of $\operatorname{Pic}(X)$ with intersection matrix $\Gamma_{2,c}$. If $c=0$, then $D_{1}$ and $D_{2}$ are both classes of irreducible curves of genus 1 and $D_{i}$ is a bisection for the genus 1 fibration given by $D_{j}$, for $i,j\in\{1,2\}$ and $i\neq j$. If $c=1$, then one can assume that $D_{1}$ is the class of a curve of genus 0. Since $(D_{2}-D_{1})^{2}=-2$ and $(D_{2}-D_{1})D_{1}>0$, $(D_{2}-D_{1})$ is an effective class, which corresponds to a smooth rational curve on the surface or to a reducible curve with at least one component which is a smooth rational curve. Since there are exactly two classes of self-intersection $-2$ in $\Gamma_{2,c}$ ($D_{2}-D_{1}$ and $D_{1}-D_{2}$) and only one of them is effective, we conclude that $(D_{2}-D_{1})$ is a rational curve. Since $(D_{2}-D_{1})\cdot D_{1}=2$, it is a rational bisection of the genus 1 fibration induced by $D_{1}$. ∎
@@
Keeping the notation introduced in the proof of Corollary 3.3, the first part follows by considering $\{D_{2},D_{2}-D_{1}\}$ as basis of the Picard group.
@@
    the result described on the genus 1 fibration induced by the pencil of the lines $\mathcal{P}$ remains true on specializations of $C_{6}$ as long as the point $p$ remains a simple node for $C_{6}$; this is equivalent to requiring that $B=\varphi_{|D_{2}|}^{-1}(p)$ is an irreducible smooth rational curve. So even under specializations of $X$ such that $B$ remains a smooth rational curve, $B$ corresponds to a smooth rational bisection and generically it is ramified on smooth fibers.
@@ 3.2. Elliptic K3 surfaces of Picard number 3
  • $\mbox{rank}(\operatorname{MW}(\pi))=1$ if and only if $d>1$ if and only if $\pi$ has no reducible fiber. In this case $\operatorname{MW}(\pi)$ is generated by a section $S_{1}$ that intersects the zero section in $d-2$ points.
@@
The result was already presented in [Gar17, Sec. 4.3], but for the reader’s convenience we recap the key ideas. Since $X$ admits an elliptic fibration, $U$ has to be primitively embedded in its Picard group, see Proposition 2.12. Since the Picard number is 3, and the Picard lattice has signature $(1,2)$, the orthogonal complement of $U$ in the Picard group is $\langle-2d\rangle$ for a certain positive $d$, and $\operatorname{Pic}(X)$ is an overlattice of finite index of $U\oplus\langle-2d\rangle$. But there are no even lattices which are overlattices of finite non-trivial index of $U\oplus\langle-2d\rangle$ in which $\langle-2d\rangle$ is primitive, so $\operatorname{Pic}(X)\simeq U\oplus\langle-2d\rangle$. This argument holds for every elliptic fibration on $X$, i.e. every elliptic fibration on $X$ is associated to an embedding of $U$ in $\operatorname{Pic}(X)$, whose orthogonal complement is $\langle-2d\rangle$. So all the properties that depend only on the intersection properties of classes in $\operatorname{Pic}(X)$ are common to all elliptic fibrations on $X$. We denote by $u_{1},u_{2},b_{3}$ the basis of $\operatorname{Pic}(X)$ on which the bilinear form is $U\oplus\langle-2d\rangle$, and we fix $F:=u_{1}$ and $\mathcal{O}=u_{2}-u_{1}$. By Lemma 2.13, the torsion part of $\operatorname{MW}(\pi)$ is trivial.
If $d=1$, then the vector $b_{3}$ is a $(-2)$-class and hence (up to a sign) the class of an irreducible component of a reducible fiber that does not intersect the zero section. By the Shioda–Tate formula, the presence of this reducible fibers is equivalent to $\operatorname{MW}(\pi)=\{1\}$.
If $d>1$, then there are no $(-2)$-classes orthogonal to $U$ and so $\pi$ has no reducible fibers. Again by the Shioda–Tate formula, this is equivalent to $\operatorname{MW}(\pi)\simeq\mathbb{Z}$. The classes of the sections of $\pi$ are
@@
and the isomorphism $\operatorname{MW}(\pi)\simeq\mathbb{Z}$ maps $\mathcal{O}$ to 00 and $S_{n}$ to $n$. In particular, if $\operatorname{Pic}(X)\simeq U\oplus\langle-2d\rangle$ and $d>1$, then $X$ admits an elliptic fibration without reducible fibers such that the generator $S_{1}$ of $\operatorname{MW}(\pi)$ intersects the zero section in $d-2$, i.e. $S_{1}\mathcal{O}=d-2$. Observe that $S_{1}\mathcal{O}=S_{-1}\mathcal{O}$ and more in general $S_{n}\mathcal{O}=S_{-n}\mathcal{O}$.∎
Recalling Notation 3.1, by the previous discussion, if $d>1$ the 17-dimensional family $\mathcal{L}_{d}$ is the family of K3 surfaces that admit an elliptic fibration $\pi$ without reducible fibers with $\mbox{rank}(\operatorname{MW}(\pi))=1$ such that the generator of $\operatorname{MW}(\pi)$ intersects the zero section in $d-2$ points.
@@
As explained in the proof of Lemma 3.8, there could be more than one elliptic fibration on a K3 surface $X\in\mathcal{L}_{d}$ for a determined values of $d$. Nevertheless if $X$ is general in its family, in particular if $\rho(X)=3$, then all these elliptic fibrations admit the same configurations of reducible fibers, their Mordell–Weil groups are isomorphic and the generator of the Mordell–Weil group intersects the zero section in the same number of points (briefly, their Mordell–Weil lattices are the same).

### v2

@@ 1. Introduction
Let $X$ be a K3 surface defined over a number field such that $\operatorname{Pic}(X)=U\oplus\langle-2d\rangle$, for $d\in\mathbb{N}$, or equivalently a K3 surface defined over $k$ admitting an elliptic fibration (possibly over a field extension of $k$) and with Picard number 3. Assume that $d\neq 2,3$. Then each elliptic fibration on $X$ has the potential rank jump property and, moreover, if $d\equiv 3\mod 4$, then it has the rank jump property over its field of definition.
@@
    The generic member in the 18-dimensional family of U(2)U(2)–polarized K3 surfaces admits a saliently ramified smooth bisection of genus 1 (Proposition 3.6).
@@ 2.1. Genus 1 fibrations and Elliptic surfaces
The set of $k$-sections of $\pi$ is an abelian group which we denote by $\operatorname{MW}(X,\pi,k)$. We often omit $k$, if the field of definition is clear from the context. The fibration $\pi:X\to C$ is called trivial if its generic fiber is isomorphic to $E\otimes_{k}k(C)$ for some elliptic curve $E$ over $k$.
@@ 2.1.1. Multisections of genus 1 fibrations
Given a genus 1 fibration $\pi:X\rightarrow C$, a geometrically integral curve $M\subset X$ is called a multisection of $\pi$ if $\pi|_{M}:M\rightarrow C$ is finite and of degree larger than 1.
@@
Our definition is slightly different from that in [BT00, Def. 3.2.]. We assume that the degree of the restriction of the fibration is larger than 1, while [BT00] assumes that it is non-zero. So, in our definition, sections are not multisections. This modification is natural in our settings. Indeed, we are interested in the ramification of the restriction of the fibration of the multisection (Def. 2.5) which clearly only makes sense for maps of degree at least 2.
@@ 2.2. Rank jump and base change
Let $M\subset X$ be a multisection of $\pi$ defined over $l/k$. Let $X_{M}$ be the relatively minimal elliptic surface obtained by resolving the singularities of $X\times_{C}M$. The projection onto $M$ induces an elliptic fibration $\pi_{M}:X_{M}\rightarrow M$ and we have an injection $\alpha:\mathrm{MW}(X,\pi,l)\hookrightarrow\mathrm{MW}(X_{M},\pi_{M},l)$ yielding
@@
Let $\iota:M\rightarrow X$ be the inclusion map and $\mathrm{id:}M\rightarrow M$, the identity map. Then $(\iota,\mathrm{id}):M\rightarrow X\times_{C}M$ induces a section $\sigma_{M}$ of $\pi_{M}$, namely the strict transform of $(\iota,\mathrm{id})$ under the birational map $X_{M}\rightarrow X\times_{C}M$. If the latter is linearly independent of the sections in the image of $\alpha$ then the inequality in (2.1) is strict and, in particular, an application of Silverman’s Specialization Theorem ([Sil83, Theorem C]) implies that $\mathrm{rank}X_{t}(l)>\mathrm{rank}\,\operatorname{MW}(X,\pi,l)$ for all but finitely many $t$ in the image of the $l$-points of the map $\pi|_{M}:M\rightarrow C$.
@@
Let $\pi:X\rightarrow C$ be an elliptic fibration and $M$ a saliently ramified multisection defined over a field $l$. Let $X_{M}$ and $\pi_{M}:X_{M}\rightarrow M$ be as above. Then the section $\sigma_{M}$ induced on $X_{M}$ by $(\iota,\mathrm{id}):M\rightarrow X\times_{C}M$ is linearly independent of the pull-back of the sections of $\pi$ in $\operatorname{MW}(X_{M},\pi_{M},l)$.
@@
$\mathcal{R}(X,\pi,l)\text{ is infinite.}$
@@ 2.3. Lattice polarized K3 surfaces and genus 1 fibrations
A K3 surface $X$ admits an elliptic fibration if, and only if, the lattice $U\simeq\left[\begin{array}[]{ll}0&1\\ 1&0\end{array}\right]$ is primitively embedded in $\operatorname{Pic}(X)$. In particular, the family of K3 surfaces admitting an elliptic fibration is 18-dimensional and coincides with the family of $U$-polarized K3 surfaces.
@@
This result is well known, cf. [Ko, Lemma 2.1]. We briefly recall the idea of the proof. If a K3 surface admits an elliptic fibration, then there exists a smooth genus 1 curve, that is, a fiber of the fibration. We denote it by $F$ and observe that $F^{2}=0$ (by adjuction). Moreover, there exists a rational curve that is a section, denoted by $\mathcal{O}$. Then $\mathcal{O}^{2}=-2$ (by adjuction) and $F\mathcal{O}=1$ (since $\mathcal{O}$ is a section). Therefore, the lattice $U\simeq\langle F,F+\mathcal{O}\rangle$ is a sublattice of $\operatorname{Pic}(X)$. Since $U$ is a unimodular lattice, it does not have overlattices, so the embedding $\langle F,F+\mathcal{O}\rangle\subset\operatorname{Pic}(X)$ is primitive. The converse is more delicate; indeed, it is not true that each class with self-intersection 0 is necessarily the class of a smooth irreducible curve (and hence of a fiber of a genus 1 fibration). Nevertheless, given a primitive class in $\operatorname{Pic}(X)$ with self-intersection 0, it is true that up to reflections induced by smooth rational curves on the surface, one can assume that it is the class of a smooth irreducible curve, which necessarily has genus 1 by adjuction. So, let $U$ be primitively embedded in $\operatorname{Pic}(X)$. Considering (potentially) reflections in 2-2-curves, one can assume that one of the classes with self-intersection 0 is the class of a smooth irreducible curve of genus 1, say $F$, and then $|F|$ defines a fibration onto $\mathbb{P}(H^{0}(X,\mathcal{O}_{X}(F))^{\vee})=\mathbb{P}^{1}$. Denoted $\{F,v\}$ the base of $U$, with $F$ as above, vFv-F is a 2-2-class with positive intersection with $F$, i.e., it is an effective 2-2-class. Moreover, it intersects the fiber in one point and therefore it is either an irreducible curve, and in particular a section, or the union of several curves, among which there is a section. In both the cases the fibration admits a section.∎
@@
If a K3 surface contains two distinct genus 1 fibrations, then the class of the fiber of the first one, say $F_{1}$, and the one of the second one, say $F_{2}$, have a positive intersection $F_{1}F_{2}=m>1$. The intersection has to be larger than 1 since the base of a genus 1 fibration on a K3 surface is rational and so $F_{2}$ cannot be a section. In particular, $\langle F_{1},F_{2}\rangle\simeq U(m)$, with $m>1$, is primitively embedded in the Picard group of the surface. The generic K3 surface $X$ such that $\operatorname{Pic}(X)\simeq U(m)$ admits two genus 1 fibrations whose classes of fibers meet in $m$ points. Therefore the K3 surfaces that admit more than one genus 1 fibration are contained in the families of $U(m)$-polarized K3 surfaces for a certain $m>1$.
@@ 2.3.3. K3 surfaces with an elliptic fibration and low Picard number
Let $(X,\pi,\mathbb{P}^{1})$ be an elliptic fibration on the K3 surface $X$. Then there are no multiple fibers of $\pi$, due to the presence of a section, nevertheless there can be reducible fibers which contain some non-reduced components. These are the fibers of type $_{n}^{*}, II$$^{*}, II$$II^{*}$ or $IV^{*}$ in the Kodaira classification of the singular fibers, see [Mir89] for notation. Each of these fibers has at least 5 components, and in particular at least 4 components which do not intersect the zero section. These components are independent in $\operatorname{Pic}(X)$, and they are also independent with respect to the classes $F$ and $\mathcal{O}$ of the fiber and of the zero section of $\pi$, respectively. Hence, the presence of a fiber with non-reduced components forces the Picard number of the surface to be at least 6.
@@
Let $X$ be a K3 surface with $\rho(X)<6$. Any elliptic fibration $\pi:X\rightarrow\mathbb{P}^{1}$ admits neither fibers with non-reduced components nor torsion sections.
@@ 3. Bisections of genus 1 and elliptic fibrations
For any other $d$, any elliptic fibration on $X$ has no reducible fibers, the Mordell–Weil group of any elliptic fibration on $X$ is isomorphic to $\mathbb{Z}$ and admits a generator that intersects the zero section in d2d-2 points.
@@ 3.1. Bisections of genus 1 fibrations
Let $X$ be a K3 surface that admits a genus 1 fibration $\pi:X\rightarrow\mathbb{P}^{1}$ without a section. The class of the fiber $F$ has trivial self-intersection, and there exists an ample class $H$ with $H^{2}=d>0$ among the generators of $\operatorname{Pic}(X)$, whose rank is forced to be at least 2. In other words, a general K3 surface $X$ with a genus 1 fibration and Picard number 2 has $\operatorname{Pic}(X)\simeq\Gamma_{b,c}$, where $\Gamma_{b,c}$ is the indefinite lattice of rank 2 defined as
@@
where one can assume $0\leq c<b$, see [vG05, Remark 4.2]. Since $\pi$ does not admits a section, the lattice $\Gamma_{b,c}$ is not isometric to $U$. Since $U$ is the unique hyperbolic even unimodular lattice of rank 2, $\det(\Gamma_{b,c})\neq-1$ and so $b\neq 1$.
@@
If $X$ admits a genus 1 fibration with a bisection and $\rho(X)=2$, then $\operatorname{Pic}(X)=\Gamma_{2,c}$ and by [vG05, Proposition 3.7], we can assume that $c=0$ or $c=1$. Let $D_{1},D_{2}$ be two generators of $\operatorname{Pic}(X)$ with intersection matrix $\Gamma_{2,c}$. If $c=0$, then $D_{1}$ and $D_{2}$ are both classes of irreducible curves of genus 1 and $D_{i}$ is a bisection for the genus 1 fibration given by $D_{j}$, for $i,j\in\{1,2\}$ and $i\neq j$. If $c=1$, then one can assume that $D_{1}$ is the class of a curve of genus 1. Since $(D_{2}-D_{1})^{2}=-2$ and $(D_{2}-D_{1})D_{1}>0$, $(D_{2}-D_{1})$ is an effective class, which corresponds to a smooth rational curve on the surface or to a reducible curve with at least one component which is a smooth rational curve. Since there are exactly two classes of self-intersection 2-2 in $\Gamma_{2,1}$ ($D_{2}-D_{1}$ and $D_{1}-D_{2}$) and only one of them is effective, we conclude that $(D_{2}-D_{1})$ is a smooth rational curve. Since $(D_{2}-D_{1})\cdot D_{1}=2$, it is a rational bisection of the genus 1 fibration induced by $D_{1}$.
To show that the two possibilities are mutually exclusive, we prove that $\pi$ admits a rational (resp. genus 1) bisection if and only if $\operatorname{Pic}(X)\simeq\Gamma_{2,1}$ (resp. $\operatorname{Pic}(X)\simeq\Gamma_{2,0})$). We already prove the only if condition. So now assume that $\pi$ admits the required bisection $B$. Then $\operatorname{Pic}(X)$ contains the class $F$, of the fiber of $\pi$, and of $B$. Moreover, $F^{2}=0$ and $FB=2$ and $B^{2}=-2$ (resp. $B^{2}=0$). Since there are no non trivial overlattices of finte index of $\langle F,B\rangle$ which could appear as Picard group of a K3 surface, it follows that $\pi$ admits a rational bisection if and only if $\operatorname{Pic}(X)\simeq\langle F,B\rangle\simeq\left[\begin{array}[]{cc}0&2\\ 2&-2\end{array}\right]\simeq\Gamma_{2,1}$, where the isometry between the last two lattices was already proven (it suffices to compute the intersection form on $\langle F,2F+B\rangle$). Similarly, $\pi$ admits a genus 1 bisection if and only if $\operatorname{Pic}(X)\simeq\langle F,B\rangle\simeq\Gamma_{2,0}$. Since $\Gamma_{2,0}$ does not contains classes with self intersect 2 and $\Gamma_{2,1}$ does, they can not be isometric. ∎
@@
Keeping the notation introduced in the proof of Corollary 3.3, the first part follows by considering $\{D_{2},D_{2}-D_{1}\}$ as base of the Picard group.
@@
    the result described on the genus 1 fibration induced by the pencil of the lines $\mathcal{P}$ remains true on certain specializations of $C_{6}$. In particular, it remains true if the specialization $C_{6}^{\prime}$ of $C_{6}$ is such that: the double cover of $\mathbb{P}^{2}$ branched on $C_{6}^{\prime}$ is still a K3 surface $X$ and the point $p$ remains a simple node for $C_{6}^{\prime}$; the latter condition guarantees that the K3 surface $X$ admits a genus 1 fibration induced by $\mathcal{P}$ and that $B=\varphi_{|D_{2}|}^{-1}(p)$ is an irreducible smooth rational curve. One automatically obtains that $B$ is a bisection of the genus 1 fibration. So, even under these specializations, $B$ is a smooth rational bisection of a genus 1 fibration on a K3 surface and it is generically ramified on smooth fibers.
@@ 3.2. Elliptic K3 surfaces of Picard number 3
  • $\mbox{rank}(\operatorname{MW}(\pi))=1$ if and only if $d>1$ if and only if $\pi$ has no reducible fiber. In this case $\operatorname{MW}(\pi)$ is generated by a section $S_{1}$ that intersects the zero section in d2d-2 points.
@@
The result was already presented in [Gar17, Sec. 4.3], but for the reader’s convenience we recap the key ideas. Since $X$ admits an elliptic fibration, $U$ has to be primitively embedded in its Picard group, see Proposition 2.12. Since the Picard number is 3, and the Picard lattice has signature $(1,2)$, the orthogonal complement of $U$ in the Picard group is $\langle-2d\rangle$ for a certain positive $d$, and $\operatorname{Pic}(X)$ is an overlattice of finite index of $U\oplus\langle-2d\rangle$. But there are no even lattices which are overlattices of finite non-trivial index of $U\oplus\langle-2d\rangle$ in which $\langle-2d\rangle$ is primitive, so $\operatorname{Pic}(X)\simeq U\oplus\langle-2d\rangle$. This argument holds for every elliptic fibration on $X$, i.e. every elliptic fibration on $X$ is associated to an embedding of $U$ in $\operatorname{Pic}(X)$, whose orthogonal complement is $\langle-2d\rangle$. So all the geometric properties that depend only on the intersection numbers of classes in $\operatorname{Pic}(X)$ are common to all elliptic fibrations on $X$. We denote by $u_{1},u_{2},b_{3}$ the base of $\operatorname{Pic}(X)$ on which the bilinear form is $U\oplus\langle-2d\rangle$, and we fix $F:=u_{1}$ and $\mathcal{O}=u_{2}-u_{1}$. By Lemma 2.13, the torsion part of $\operatorname{MW}(\pi)$ is trivial.
If $d=1$, then the vector $b_{3}$ is a (2)(-2)-class and hence (up to a sign) the class of an irreducible component of a reducible fiber that does not intersect the zero section. By the Shioda–Tate formula, the presence of this reducible fiber is equivalent to $\operatorname{MW}(\pi)=\{1\}$.
If $d>1$, then there are no (2)(-2)-classes orthogonal to $U$ and so $\pi$ has no reducible fibers. Again by the Shioda–Tate formula, this is equivalent to $\operatorname{MW}(\pi)\simeq\mathbb{Z}$. The classes of the sections of $\pi$ are
@@
and the isomorphism $\operatorname{MW}(\pi)\simeq\mathbb{Z}$ maps $\mathcal{O}$ to 0 and $S_{n}$ to $n$. In particular, if $\operatorname{Pic}(X)\simeq U\oplus\langle-2d\rangle$ and $d>1$, then $X$ admits an elliptic fibration without reducible fibers such that the generator $S_{1}$ of $\operatorname{MW}(\pi)$ intersects the zero section in d2d-2, i.e. $S_{1}\mathcal{O}=d-2$. Observe that $S_{1}\mathcal{O}=S_{-1}\mathcal{O}$ and more in general $S_{n}\mathcal{O}=S_{-n}\mathcal{O}$.∎
Recalling Notation 3.1, by the previous discussion, if $d>1$ the 17-dimensional family $\mathcal{L}_{d}$ is the family whose general members are K3 surfaces that admit an elliptic fibration $\pi$ without reducible fibers with $\mbox{rank}(\operatorname{MW}(\pi))=1$ such that the generator of $\operatorname{MW}(\pi)$ intersects the zero section in d2d-2 points.
@@
As explained in the proof of Lemma 3.8, there could be more than one elliptic fibration on a K3 surface $X\in\mathcal{L}_{d}$ for a determined values of $d$. Nevertheless, if $X$ is general in its family, in particular if $\rho(X)=3$, then all these elliptic fibrations admit the same configurations of reducible fibers, their Mordell–Weil groups are isomorphic and the generator of the Mordell–Weil group intersects the zero section in the same number of points (briefly, their Mordell–Weil lattices are isometric).